OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all possible rational zeros of the function. f(x) = 3x^3 + 39x^2 + 39x + 27

hartnn (hartnn):

do u know rational zero theorem ??

OpenStudy (anonymous):

not really, ive been trying to figure this out... but my textbook isnt very clear

hartnn (hartnn):

then this reference might help, see solved example: http://en.wikipedia.org/wiki/Rational_root_theorem

hartnn (hartnn):

if u still don't understand, i'll explain.

OpenStudy (anonymous):

yeah.. that would be helpful

hartnn (hartnn):

ok list the factors of 27 and list the factors of 3

OpenStudy (anonymous):

1,3,9,27 and 1, 3

hartnn (hartnn):

yup, now comparing your equation with ax^2+bx+cx+d u get d=27,a=3 right According to that theorem , the possible roots are given by all possible ratios in \(\huge \frac{d}{a}=\frac{27}{3}=\frac{1,3,9,27}{1,3}\) so what all possibilities u have??

OpenStudy (anonymous):

you have... +/-.... 1, 3/1, 9/1, 27/1, 1/3, 9/3, 27/3 i think

hartnn (hartnn):

yup, there is \(\huge \pm\) sorry , i forgot that....so all those possibilities with negative sign also..... ok ?

OpenStudy (anonymous):

mmk, thank you so much! it makes way more sense now

hartnn (hartnn):

just note that u wrote 3/1 twice .... 3/1 and 9/3 ....

hartnn (hartnn):

and 9/1 twice .... 9/1 and 27/3 ....